# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

### v2

@@ Abstract
We then introduce and analyze new multivariate problems in which ${\mathbf{x}}_{i},\mathbf{y}_{i}\in\mathbb{R}^{d}$ with variational penalties that depend on $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. The norms we consider are $\ell_{2}$ and $\ell_{\infty}$ which promote group sparsity. We also derive a novel lattice-based procedure for subgradient following of variational penalties characterized through the output of arbitrary convolutional filters. This paradigm yields efficient solvers for high-order filtering problems of temporal sequences in which sparse discrete derivatives such as acceleration and jerk are desirable.
@@ 1 Sub-Introduction
We next generalize the Untangled Lariat to multivariate settings in which each element of the input sequence is a vector. We thus seek a solution ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ where ${\mathbf{x}}_{i}\in\mathbb{R}^{d}$ for all $i\in[n]$. In order to obtain variational sparsity in the multivariate case we replace the absolute value $|x_{i+1}-x_{i}|$ with norm penalties over the differences $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. To start, we derive a closed-form for the squared $2$-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}$. We use the closed form solution to build surrogate functions for the $2$-norm itself, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$, and the infinity-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}$. The result is an iterative yet efficient solver for the problem,
@@
Last, but not least, we introduce a construction from signal processing into our setting by penalizing the convolution of ${\mathbf{x}}$ with a predefined convolution filter $\boldsymbol{\tau}$. Concretely, we provide an efficient procedure for finding the optimum of,
@@ 3.1 Revisiting the Fused Lasso
For convenience we define $\mathsf{a}_{i}^{\prime}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. In order to evaluate,
$S_{i}=\{(\mathsf{x},\mathsf{a}):\mathsf{x}-\mathsf{x}^{\prime}\in\partial g_{i}^{*}(\mathsf{a})\;;\;\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})\}$
we need to analyze $\Delta\mathsf{x}=\mathsf{x}-\mathsf{x}^{\prime}$ for the admissible set of values of $\mathsf{a}$ under the map $g_{i}^{*}$ which is $[-\lambda_{i},\lambda_{i}]$. Since $\mathsf{a}_{i+1}(\cdot)$ is continuous and non-decreasing, it implies that $\mathsf{a}_{i}^{\prime}(.)$ is strictly-increasing and bijective. Therefore, there exists two unique boundary values $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=-\lambda_{i}$ and $z_{i}^{+}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{+})=+\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})$ such that $-\lambda_{i}<\mathsf{a}<\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=\{0\}$. Thus, $\mathsf{x}=\mathsf{x}^{\prime}$ and both are in $(z_{i}^{-},z_{i}^{+})$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\in(z_{i}^{-},z_{i}^{+})$ yields,
$S_{i}\cap\{\mathsf{x}:z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}$ $=$ $\{(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x})):z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}$
@@
We last need to address the cases $\mathsf{a}=\pm\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(z_{i}^{+})=\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=[0,\infty)$. Thus, for any $\mathsf{x}\geq z_{i}^{+}$ we have $\mathsf{x}-z_{i}^{+}\in\partial g^{*}_{i}(\lambda_{i})$ and therefore $(\mathsf{x},\lambda_{i})\in S_{i}$. Analogously, for $\mathsf{a}=-\lambda_{i}$ which implies that for $\mathsf{x}\leq z_{i}^{-}$, it holds that $\mathsf{x}-z_{i}^{-}\in\partial g^{*}_{i}(-\lambda_{i})\Rightarrow(\mathsf{x},-\lambda_{i})\in S_{i}$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\geq z_{i}^{+}$ and $\mathsf{x}\leq z_{i}^{-}$ yields,
@@
Since $\mathsf{a}_{i}^{\prime}(\cdot)$ is increasing, we have for $\mathsf{x}\geq z_{i}^{+}\Rightarrow\mathsf{a}_{i}^{\prime}(\mathsf{x})\geq\lambda_{i}$ and $\mathsf{x}\leq z_{i}^{-}\Rightarrow\mathsf{a}_{i}^{\prime}(\mathsf{x})\leq-\lambda_{i}$. Therefore, we established that
@@ 3.4 Isotonic Regression
where $[z]_{+}=\max(z,0)$. The derivation follows the same lines as above as there exists a unique boundary value $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=0$. The resulting description of $S_{i}$ is obtained from the union of,
@@ 3.5 Fused Barriers
Albeit being slightly more involved, we next show that there exists a succinct representation of $S_{i}$ in the form of a piecewise linear map, $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$. First, let us reuse the intermediate function $\mathsf{a}_{i}^{\prime}$ from the derivation of the fused lasso algorithm, $\mathsf{a}_{i}^{\prime}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. We denote by $\kappa_{\lambda}(\cdot)$ the Kvetsh operator (from Yiddish, literally “to squeeze, shrink”), colloquially called shrinkage,
@@
$\mathsf{a}_{i}^{\prime}(\mathsf{x})$ $:=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$ (Add Coordinate)
$z_{i}$ $:=z\,\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z)=0$ (Zero Crossing)
$\mathsf{a}_{i}(\mathsf{x})$ $:=\mathsf{a}_{i}^{\prime}\big{(}z_{i}+\kappa_{\lambda_{i}}(\mathsf{x}-z_{i})\big{)}$ (Kvetsh)
@@
We now provide the details of the derivation of $\mathsf{a}_{i}$. Since $z_{i}$ is the locus of the zero of $\mathsf{a}^{\prime}_{i}$ then for $\mathsf{x}^{\prime}>z_{i}$ we have $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})>0\Rightarrow\partial g_{i}^{*}(\mathsf{a})=\lambda_{i}$ based on Eq. (9). Thus, for any $\mathsf{x}=\mathsf{x}^{\prime}+\lambda_{i}>z_{i}+\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}-\lambda_{i}))\in S_{i}$. Analogously, for $\mathsf{x}<z_{i}-\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\in S_{i}$. Last, for $(z_{i},\mathsf{a}^{\prime}_{i}(z_{i}))=(z_{i},0)$ we have $\partial g_{i}^{*}(0)=[-\lambda_{i},\lambda_{i}]$. Therefore, for any $\mathsf{x}$ such that $z_{i}-\lambda_{i}\leq\mathsf{x}\leq z_{i}+\lambda_{i}$ it holds that $\mathsf{x}-z_{i}\in\partial g_{i}^{*}(0)\Rightarrow(\mathsf{x},0)\in S_{i}$. The resulting $S_{i}$ is obtained from the union of,
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}<z_{i}-\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i}^{\prime}(\mathsf{x}-\lambda_{i}))\,:\,\mathsf{x}<z_{i}-\lambda_{i}\}$
@@
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}~{}~{}.$
@@ 5.2 Multivariate Regression With $\infty$-norm Variational Penalty
6 High-Order Variational Penalties
Analogously, we define a sequence of intermediate functions,
We next show that the following recursive form holds,
Now, given $S_{i+1}$ we perform the following expansion,
$\min_{{\mathbf{x}}_{1:k}}f_{k}({\mathbf{x}})+\sum_{i=1}^{k}h_{i}(x_{i})~{}~{}.$
The subgradient conditions imply that
[…]
Second-order Variational Penalties.
$S_{i}=\{\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})\;;\;\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})\}~{}\mbox{ and }~{}S_{i+\frac{1}{2}}=\{\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}~{}~{}.$
$\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})=2\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})-\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})+\mathsf{x}_{i+1}-y_{i+1}~{}\mbox{ and }~{}\mathsf{a}_{i+1}^{\prime}(\cdot)=\mathsf{a}_{i+1}(\cdot)~{}~{}.$
$2\mathsf{x}_{i}-\mathsf{x}_{i-1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1}))~{}~{}.$
where $z_{i+1}^{\pm}(\mathsf{x}_{i})$ denotes the boundary values such that $\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},z_{i+1}^{\pm}(\mathsf{x}_{i}))=\pm 1$. Using this form of $T$ back in $\mathsf{x}_{i+1}$, we obtain
$\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))$
and
$\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))~{}~{}.$
@@ Appendix A Improved Barrier Lariat using Skiplists
We store each $\mathsf{a}_{i}(\cdot)$ as a linked-list data structure of implicit change points (vertices) and explicit linear intervals connecting two consecutive vertices. A vertex $(v_{j},\mathsf{a}_{i}(v_{j}))=(v_{j},\alpha_{j})$ represents a point where $\mathsf{a}_{i}(\cdot)$ changes slope. An edge $(d_{j},s_{j})$ stores the length and slope of the line segment between $v_{j}$ and $v_{j+1}$, thus we have,
$d_{j}=v_{j+1}-v_{j}\,;\,s_{j}=(\alpha_{j+1}-\alpha_{j})/d_{j}~{}~{}.$
This representation allows lazy evaluations in a similar fashion to the fused lasso. The Kvetsh step does not change the lengths of intervals except for the new segment encapsulating the zero crossing. Since the length of all other line segments remain intact this edge-based representation facilitates an efficient implementation.

### v4

@@ Abstract
We then derive a novel lattice-based procedure for subgradient following of variational penalties characterized through the output of arbitrary convolutional filters. This paradigm yields efficient solvers for high-order filtering problems of temporal sequences in which sparse discrete derivatives such as acceleration and jerk are desirable. We also introduce and analyze new multivariate problems in which ${\mathbf{x}}_{i},\mathbf{y}_{i}\in\mathbb{R}^{d}$ with variational penalties that depend on $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. The norms we consider are $\ell_{2}$ and $\ell_{\infty}$ which promote group sparsity.
@@ 1 Sub-Introduction
We next introduce a construction from signal processing into our setting by penalizing the convolution of ${\mathbf{x}}$ with a predefined convolution filter $\boldsymbol{\tau}$. Concretely, we provide an efficient procedure for finding the optimum of,
Last but not least, we generalize the Untangled Lariat to multivariate settings in which each element of the input sequence is a vector. We thus seek a solution ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ where ${\mathbf{x}}_{i}\in\mathbb{R}^{d}$ for all $i\in[n]$. In order to obtain variational sparsity in the multivariate case we replace the absolute value $|x_{i+1}-x_{i}|$ with norm penalties over the differences $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. To start, we derive a closed-form for the squared $2$-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}$. We use the closed form solution to build surrogate functions for the $2$-norm itself, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$, and the infinity-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}$. The result is an iterative yet efficient solver for the problem,
@@ 3.1 Revisiting the Fused Lasso
For convenience we define $\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. In order to evaluate,
$S_{i}=\{(\mathsf{x},\mathsf{a}):\mathsf{x}-\mathsf{x}^{\prime}\in\partial g_{i}^{*}(\mathsf{a})\;;\;\mathsf{a}=\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}^{\prime})\}$
we need to analyze $\Delta\mathsf{x}=\mathsf{x}-\mathsf{x}^{\prime}$ for the admissible set of values of $\mathsf{a}$ under the map $g_{i}^{*}$ which is $[-\lambda_{i},\lambda_{i}]$. Since $\mathsf{a}_{i+1}(\cdot)$ is continuous and non-decreasing, it implies that $\mathsf{a}_{i+\frac{1}{2}}(.)$ is strictly-increasing and bijective. Therefore, there exists two unique boundary values $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i+\frac{1}{2}}(z_{i}^{-})=-\lambda_{i}$ and $z_{i}^{+}\;\text{s.t.}\;\mathsf{a}_{i+\frac{1}{2}}(z_{i}^{+})=+\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}^{\prime})$ such that $-\lambda_{i}<\mathsf{a}<\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=\{0\}$. Thus, $\mathsf{x}=\mathsf{x}^{\prime}$ and both are in $(z_{i}^{-},z_{i}^{+})$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\in(z_{i}^{-},z_{i}^{+})$ yields,
$S_{i}\cap\{\mathsf{x}:z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}$ $=$ $\{(\mathsf{x},\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x})):z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}$
@@
We last need to address the cases $\mathsf{a}=\pm\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i+\frac{1}{2}}(z_{i}^{+})=\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=[0,\infty)$. Thus, for any $\mathsf{x}\geq z_{i}^{+}$ we have $\mathsf{x}-z_{i}^{+}\in\partial g^{*}_{i}(\lambda_{i})$ and therefore $(\mathsf{x},\lambda_{i})\in S_{i}$. Analogously, for $\mathsf{a}=-\lambda_{i}$ which implies that for $\mathsf{x}\leq z_{i}^{-}$, it holds that $\mathsf{x}-z_{i}^{-}\in\partial g^{*}_{i}(-\lambda_{i})\Rightarrow(\mathsf{x},-\lambda_{i})\in S_{i}$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\geq z_{i}^{+}$ and $\mathsf{x}\leq z_{i}^{-}$ yields,
@@
Since $\mathsf{a}_{i+\frac{1}{2}}(\cdot)$ is increasing, we have for $\mathsf{x}\geq z_{i}^{+}\Rightarrow\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x})\geq\lambda_{i}$ and $\mathsf{x}\leq z_{i}^{-}\Rightarrow\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x})\leq-\lambda_{i}$. Therefore, we established that
@@ 3.4 Isotonic Regression
where $[z]_{+}=\max(z,0)$. The derivation follows the same lines as above as there exists a unique boundary value $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i+\frac{1}{2}}(z_{i}^{-})=0$. The resulting description of $S_{i}$ is obtained from the union of,
@@ 3.5 Fused Barriers
Albeit being slightly more involved, we next show that there exists a succinct representation of $S_{i}$ in the form of a piecewise linear map, $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$. First, let us reuse the intermediate function $\mathsf{a}_{i+\frac{1}{2}}$ from the derivation of the fused lasso algorithm, $\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. We denote by $\kappa_{\lambda}(\cdot)$ the Kvetsh operator (from Yiddish, literally “to squeeze, shrink”), colloquially called shrinkage,
@@
$\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x})$ $:=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$ (Add Coordinate)
$z_{i}$ $:=z\,\;\text{s.t.}\;\mathsf{a}_{i+\frac{1}{2}}(z)=0$ (Zero Crossing)
$\mathsf{a}_{i}(\mathsf{x})$ $:=\mathsf{a}_{i+\frac{1}{2}}\big{(}z_{i}+\kappa_{\lambda_{i}}(\mathsf{x}-z_{i})\big{)}$ (Kvetsh)
@@
We now provide the details of the derivation of $\mathsf{a}_{i}$. Since $z_{i}$ is the locus of the zero of $\mathsf{a}^{\prime}_{i}$ then for $\mathsf{x}^{\prime}>z_{i}$ we have $\mathsf{a}=\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}^{\prime})>0\Rightarrow\partial g_{i}^{*}(\mathsf{a})=\lambda_{i}$ based on Eq. (9). Thus, for any $\mathsf{x}=\mathsf{x}^{\prime}+\lambda_{i}>z_{i}+\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}-\lambda_{i}))\in S_{i}$. Analogously, for $\mathsf{x}<z_{i}-\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}+\lambda_{i}))\in S_{i}$. Last, for $(z_{i},\mathsf{a}^{\prime}_{i}(z_{i}))=(z_{i},0)$ we have $\partial g_{i}^{*}(0)=[-\lambda_{i},\lambda_{i}]$. Therefore, for any $\mathsf{x}$ such that $z_{i}-\lambda_{i}\leq\mathsf{x}\leq z_{i}+\lambda_{i}$ it holds that $\mathsf{x}-z_{i}\in\partial g_{i}^{*}(0)\Rightarrow(\mathsf{x},0)\in S_{i}$. The resulting $S_{i}$ is obtained from the union of,
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}<z_{i}-\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}-\lambda_{i}))\,:\,\mathsf{x}<z_{i}-\lambda_{i}\}$
@@
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}+\lambda_{i}))\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}~{}~{}.$
@@
4 High-Order Variational Penalties
Analogously, we define a sequence of intermediate functions,
We next show that the following recursive form holds,
Now, given $S_{i+1}$ we perform the following expansion,
$\min_{{\mathbf{x}}_{1:k}}f_{k}({\mathbf{x}})+\sum_{i=1}^{k}h_{i}(x_{i})~{}~{}.$
The subgradient conditions imply that
[…]
Second-order Variational Penalties.
$S_{i}=\{\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})\;;\;\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})\}~{}\mbox{ and }~{}S_{i+\frac{1}{2}}=\{\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+\frac{3}{2}}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}~{}~{}.$
$\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}_{i},\mathsf{x}_{i+1})=2\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})-\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})+\mathsf{x}_{i+1}-y_{i+1}~{}\mbox{ and }~{}\mathsf{a}_{i+\frac{3}{2}}(\cdot)=\mathsf{a}_{i+1}(\cdot)~{}~{}.$
$2\mathsf{x}_{i}-\mathsf{x}_{i-1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}_{i},\mathsf{x}_{i+1}))~{}~{}.$
where $z_{i+1}^{\pm}(\mathsf{x}_{i})$ denotes the boundary values such that $\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}_{i},z_{i+1}^{\pm}(\mathsf{x}_{i}))=\pm 1$. Using this form of $T$ back in $\mathsf{x}_{i+1}$, we obtain
$\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i+\frac{1}{2}}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))$
and
$\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i+\frac{3}{2}}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))~{}~{}.$
@@ Appendix A Improved Barrier Lariat using Skiplists
We store each $\mathsf{a}_{i}(\cdot)$ as a linked-list data structure of implicit change points (vertices) and explicit linear intervals connecting two consecutive vertices. A vertex $(v_{j},\mathsf{a}_{i}(v_{j}))=(v_{j},\alpha_{j})$ represents a point where $\mathsf{a}_{i}(\cdot)$ changes slope. An edge $(d_{j},s_{j})$ stores the length and slope of the line segment between $v_{j}$ and $v_{j+1}$, thus we have, $d_{j}=v_{j+1}-v_{j}\,;\,s_{j}=(\alpha_{j+1}-\alpha_{j})/d_{j}~{}~{}.$ This representation allows lazy evaluations in a similar fashion to the fused lasso. The Kvetsh step does not change the lengths of intervals except for the new segment encapsulating the zero crossing. Since the length of all other line segments remain intact this edge-based representation facilitates an efficient implementation.